OpenStudy (errinss):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 196 A. 14 B. –14, 14 C. 98 D. –98, 98

pooja195 (pooja195):

Any ideas on where to start?

OpenStudy (errinss):

well, i know i cant divide 162 by 2 cus that wouldnt give me the right answer uhm... should i multiply 14 by 14 to see what i get?

OpenStudy (errinss):

whoops i mean 196 lol

pooja195 (pooja195):

Hmm not quite : \[\huge~\rm~\bf~x^{2}=196\] Keep in mind there is a square (^2) how would we get rid of "x^2"

OpenStudy (errinss):

uhm i dont know....

OpenStudy (errinss):

but i do know 14*14 is 196 and -14*-14 also is 196

pooja195 (pooja195):

You answers are somewhat correct, to solve this we would square root both sides \[\huge~\rm~\bf~x^{2}=196\] \[\huge~\rm~\bf \sqrt{x^2}=\sqrt{196}\] you get 14 but you see, 14^2=196 -14^2=196 cuz -14 x -14=196

OpenStudy (errinss):

\[4x^{2}=121\]

pooja195 (pooja195):

what do we do first?

OpenStudy (errinss):

divide 4 by both sides?

pooja195 (pooja195):

perfect 121/4=?

OpenStudy (errinss):

121/4=30.25

pooja195 (pooja195):

in this case since it's a decimal let's leave it as a fraction now we have \[\huge~\rm~\bf~\sqrt{x^2}=\sqrt{\frac{ 121 }{ 4}}\] solve

OpenStudy (errinss):

60.5? i think...

OpenStudy (errinss):

oh wait lol no its \[\frac{ 11 }{ 2}\]

pooja195 (pooja195):

Good turn that fraction into a decimal what do we get?

OpenStudy (errinss):

5.5?

pooja195 (pooja195):

perfect but what other answer could it be?

OpenStudy (errinss):

.18 repeating?

pooja195 (pooja195):

nope it's 5.5 but what's the opposite of 5.5 (look at the previous problem) I said: you get 14 but you see, 14^2=196 -14^2=196 cuz -14 x -14=196 So what would our answers be here?

OpenStudy (errinss):

-5.5

pooja195 (pooja195):

and?

OpenStudy (errinss):

5.5

pooja195 (pooja195):

perfect :-)

OpenStudy (errinss):

another?

pooja195 (pooja195):

new post :P